In a manufacturing process a random sample of 9 bolts manufactured has a mean length of 3 inches with a variance of .09. What is the 90% confidence interval for the true mean length of the bolt?

A. 2.8355 to 3.1645

B. 2.5065 to 3.4935

C. 2.8140 to 3.1860

D. 2.4420 to 3.5580

E. 2.9442 to 3.0558
Answer #1

Solution :

Given that,

t /2,df = 1.860

Margin of error = E = t/2,df * (s /n)

= 1.860 * (0.3 / 9)

Margin of error = E = 0.186

The 90% confidence interval estimate of the population mean is,

- E < < + E

3 - 0.186 < < 3 + 0.186

2.8140 < < 3.1860

(2.8140 , 3.1860)

option C. is correct